# Some Optimizers are More Equal: Understanding the Role of Optimizers in Group Fairness

### v1

@@ 3.3 Group fairness analysis
For SGD, the difference between the parameter updates across subgroups can be quantified as $||\nabla\mathcal{L}_{0}-\nabla\mathcal{L}_{1}||$. However, as the RMSProp scales the updates using exponentially decaying average, computing the difference is more complex. Let’s compute the expected value of $v_{k}$ defined in Eq. 2:
@@
where $Z$ is a random variable denoting the subgroup. Finally, assuming that (a) $\Theta_{0}^{2}$ and $\Theta_{1}^{2}$ are larger than $\mu_{0}^{2}$ and $\mu_{1}^{2}$, as shown experimentally in prior work [NEURIPS2022_32ac7101], and (b) there is bias in the training set (i.e., $p_{0}p_{1}$ is small), the variance $\operatorname{var}(v_{k_{j}})$ can be computed as $\mathcal{O}((1-\gamma)(\Theta_{0_{j}}^{4}+\Theta_{1_{j}}^{4}))$. For a large $\gamma$, typical in practical training, $var(v_{k_{j}})$ becomes negligible, and thus $v_{k}\rightarrow E[v_{k}]$. Using this, we can approximate the difference between the parameter updates across subgroups for RMSProp as $||D(\nabla\mathcal{L}_{0}-\nabla\mathcal{L}_{1})||$, where $D$ is a diagonal matrix, whose $jj$th element can be computed as: $D_{jj}=\frac{1}{\sqrt{p_{0}\left(\mu_{0_{j}}^{2}+\Theta_{0_{j}}^{2}\right)+p_{1}\left(\mu_{1_{j}}^{2}+\Theta_{1_{j}}^{2}\right)+\epsilon}}$. As $\Theta^{2}>\mu^{2}$ and in stochastic (online) training setup, we can set $\Theta^{2}\rightarrow 1$, then $D_{jj}<1$. In other words, the eigenvalues of matrix $D$ is less than 1, thus proving the theorem. ∎
Theorem 2 shows that RMSProp, unlike SGD, shrinks disparities between subgroups in the updates. In other words, RMSProp prevents large gradient disparities from dominating the training dynamics. Consequently, the final model is less likely to be pulled toward the subgroup with the larger gradients. Over many iterations, this can help prevent large performance gaps between subgroups. While this does not establish a universal fairness guarantee, it outlines why RMSProp’s dynamics yield fairer minima across subgroups. The following theorem analyzes fairness through the lens of demographic parity and establishes that, under appropriate conditions, in a single iteration, RMSProp’s worst-case increase in the gap of demographic parity has an upper-bound no greater than that of SGD.
@@
We can compute $\nabla\mathcal{F}(w_{k})$ as:
$\nabla\mathcal{F}(w_{k})=\varphi(w_{k})\cdot(\nabla\Psi(w_{k})).$ (27)
Similarity, for RMSProp, we have:
In the equation above, as diagonal elements of $D$ are strictly less than 1, we have a strict inequality. Comparing Eqs. 30 and 31, we conclude the proof. ∎
This theorem establishes that the adaptive learning rate of RMSProp, which scales updates based on past squared gradients, helps mitigate demographic parity gaps that can occur during training, whereas SGD does not offer such a mechanism.
@@ 4 Experiments
Setup. We use three public datasets that are widely used in this area: CelebA [liu2015faceattributes], FairFace [karkkainenfairface], and MS-COCO [10.1007/978-3-319-10602-1_48]. CelebA and FairFace are datasets commonly used for group fairness analysis, while we also use a general-purpose dataset, MS-COCO, for further completeness. The details about these datasets are presented in Appendix I. The training protocol, backbones, and fairness measures used in our experiments are described in Appendix F and G, respectively.
Implementation details. We use PyTorch [paszke2019pytorch] and up to four Nvidia A100 GPUs. The hyperparameter values are provided in Appendix J, which have been optimized using Weights & Biases [wandb].
@@
Regarding $F_{DPA}$, Adam consistently outperforms SGD, while RMSProp surpasses SGD in 4 out of 5 scenarios. Both optimizers achieve zero $F_{DPA}$ on MS-COCO with gender as the sensitive attribute. Interestingly, the largest discrepancy between RMSProp and SGD occurs on the FairFace dataset with age as the sensitive attribute. This aligns with our theoretical insights from Theorem 1. Specifically, FairFace exhibits extreme imbalance, with a minority-to-all ratio of approximately 0.9%, whereas CelebA with gender as the sensitive attribute has a substantially higher ratio of 42%. These findings highlight that as dataset imbalance intensifies, the gap between adaptive optimizers and SGD becomes more evident. We refer the reader to Appendix K for the results on fairness metrics for other backbones across CelebA, FairFace, and MS-COCO datasets with different sensitive attributes. The results are well-aligned with the results above.
@@
To demonstrate that our training is successful and the models are well-behaved, we evaluate the classification performance of ViT on the three datasets for facial expression recognition, multi-label classification, and gender classification, respectively. The results are presented in Table 1. The performance of the other models are presented in Appendix L. Here, we can confirm that in most cases, all three optimizers exhibit comparable accuracy and F1 scores across all tasks. For instance, on the CelebA dataset, SGD and RMSProp achieve an accuracy of 91%, while Adam attains 92%. This trend is consistent with previous findings in the prior work [maunderstanding]. As evident, the improved fairness observed with adaptive gradient algorithms cannot be solely attributed to their performance. For instance, on the CelebA dataset, the F1 score for SGD is higher than that of RMSProp, whereas on FairFace, RMSProp achieves a higher F1 score than SGD. However, in both cases, RMSProp demonstrates better fairness performance, as shown in Fig. 2. This further supports our theoretical conclusion that RMSProp is fairer than SGD, regardless of overall performance.
@@
$F_{EOD}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$F_{EOP}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-3}$
@@
To rigorously assess the significance of fairness improvement of Adam and RMSProp over SGD, we conduct the following experiment. We train ViT on CelebA dataset 10 times using SGD, RMSProp, and Adam, while recording fairness metrics for both gender and age attributes in each run. To evaluate statistical significance, we perform a paired Wilcoxon signed-rank test [benavoli2014bayesian] on the fairness metrics and report the corresponding p-values in Table 2. As shown in the table, most scenarios yield p-values in the order of 0.001, indicating a statistically significant improvement in fairness for adaptive gradient optimizers compared to SGD. The worst p-value in the table corresponds to SGD vs. RMSProp on CelebA with age as the sensitive attribute. However, even this value remains statistically significant under the conventional threshold of 0.05 [NEURIPS2024_ce7984e3, paleja2021utility].
@@ 5 Conclusion
This paper provided both theoretical and empirical evidence that the choice of optimizer can significantly affect the group fairness of resulting deep learning models. Through establishing an analytically tractable setup and analyzing it via stochastic differential equations, we showed that SGD and RMSProp can converge to different minima with regards to fairness, with RMSProp more frequently reaching fairer optima. Next, we provided two theorems that showed 1) adaptive gradient methods (e.g., RMSProp) shrink subgroup-specific gradients more effectively, thereby reducing unfair updates compared to their SGD counterparts; 2) In a single optimization step, the worst case domographic disparity of RMSProp has the upper bound given by that of SGD. We validated these theoretical findings on three diverse datasets, namely CelebA, FairFace, and MS-COCO, and on multiple tasks. Across a range of group fairness criteria, i.e, equalized odds, equal opportunity, and demographic parity, adaptive optimizers consistently led to fairer solutions. We hope our work will encourage broader adoption of adaptive optimizers in fairness-critical domains, and we envision this research as a step toward developing more equitable deep learning models.
@@ Acknowledgment
We would like to thank anonymous.
@@ Appendix B Fokker-Planck equation
Also, suppose that $Q(x)$ is a function which is 00 outside of a bounded interval. For simplicity, we use $X$ instead of $X(t)$. By Itô’s lemma, we have:
@@
[…]
@@
[…]
@@
[…]
@@ Appendix C Well-behaved NGOS
Noisy Gradient Oracle with Scale Parameter (NGOS) models stochastic gradients as $g(w)=\nabla\mathcal{L}(w)+\Theta z$, where $\Theta$ is a noise scale parameter, and $z$ is a random variable that follows a distribution with covariance matrix $\Sigma(w)$ and a mean of zero. An NGOS is well-behaved if the conditions below are satisfied [NEURIPS2022_32ac7101]:
@@
5. Noise distribution must have bounded moments. In other words, there must exist a constant $\kappa$ such that $\mathbb{E}[||z||^{2k}]^{\frac{1}{2k}}\leq\kappa(1+||w||)$.
Gaussian noise is a well-behaved NGOS. However, the noise in NGOS is not restricted to Gaussian. Particularly, if the noise distribution is not heavy-tailed, then it satisfies requirements 4 and 5. Whether or not the noise distribution in real-world applications is heavy-tailed is still an open problem. Nevertheless, several experimental results denote that the conditions described above are not too strong [NEURIPS2022_32ac7101].
@@ Appendix D Approximation quality of SDE for SGD
The dynamics of SGD can be approximated by the following SDE [li2017stochastic]:
@@
where $W_{t}$ represents an approximation of $w_{k}$ at discrete time steps $k\,\eta$ and $\Sigma$ is the covariance matrix of the underlying NGOS. The noises that make the individual paths for the above SDE and the SGD are independent processes. Hence, SGD and its SDE approximation do not share the same sample paths; rather, they are close to each other in “distribution” or “weak sense” [li2017stochastic]. To formalize this mathematically, suppose $\Xi$ be the set of functions, $f(x)$, of polynomial growth, i.e. $\exists a,b\in\mathbb{R}^{+}$ such that $|f(x)|<a(1+|x|^{b})$. Then, SDE defined in Eq. 42 is the order $r$ weak approximation of SGD if for each function $f(x)$ in $\Xi$, there exists a constant $C$ such that for all $k=1\dots N$, $|\mathbb{E}[f(W_{k\eta})]-\mathbb{E}[f(w_{k})]|<C\eta^{r}$. This is a well-known result in analyzing optimization algorithms using SDEs, first time proposed in [li2017stochastic]. Similar theorems for RMSProp and Adam are available in [NEURIPS2022_32ac7101].

### v2

@@ 3.3 Group fairness analysis
For SGD, the difference between the parameter updates across subgroups can be quantified as $\nabla\mathcal{L}_{0}-\nabla\mathcal{L}_{1}||$. However, as the RMSProp scales the updates using exponentially decaying average, computing the difference is more complex. Let’s compute the expected value of $v_{k}$ defined in Eq. 2:
@@
where $Z$ is a random variable denoting the subgroup. Finally, assuming that (a) $\Theta_{0}^{2}$ and $\Theta_{1}^{2}$ are larger than $\mu_{0}^{2}$ and $\mu_{1}^{2}$, as shown experimentally in prior work [37], and (b) there is bias in the training set (i.e., $p_{0}p_{1}$ is small), the variance $\operatorname{var}(v_{k_{j}})$ can be computed as $\mathcal{O}((1-\gamma)(\Theta_{0_{j}}^{4}+\Theta_{1_{j}}^{4}))$. For a large $\gamma$, typical in practical training, $var(v_{k_{j}})$ becomes negligible, and thus $v_{k}\rightarrow E[v_{k}]$. Using this, we can approximate the difference between the parameter updates across subgroups for RMSProp as $D(\nabla\mathcal{L}_{0}-\nabla\mathcal{L}_{1})||$, where $D$ is a diagonal matrix, whose $jj$th element can be computed as: $D_{jj}=\frac{1}{\sqrt{p_{0}\left(\mu_{0_{j}}^{2}+\Theta_{0_{j}}^{2}\right)+p_{1}\left(\mu_{1_{j}}^{2}+\Theta_{1_{j}}^{2}\right)+\epsilon}}$. As $\Theta^{2}>\mu^{2}$ and in stochastic (online) training setup, we can set $\Theta^{2}\rightarrow 1$, then $D_{jj}<1$. In other words, the eigenvalues of matrix $D$ is less than 1, thus proving the theorem. ∎
Theorem 2 shows that RMSProp, unlike SGD, shrinks disparities between subgroups in the updates. In other words, RMSProp prevents large gradient disparities from dominating the training dynamics. Consequently, the final model is less likely to be pulled toward the subgroup with the larger gradients. Over many iterations, this can help prevent large performance gaps between subgroups. While this does not establish a universal fairness guarantee, it outlines why RMSProp’s dynamics yield fairer minima across subgroups. We relax the isotropic noise assumption in Theorem 2 and generalize it to the anisotropic case in Appendix G. The following theorem analyzes fairness through the lens of demographic parity and establishes that, under appropriate conditions in a single iteration, RMSProp’s worst-case increase in the gap of demographic parity has an upper-bound no greater than that of SGD.
@@
See Appendix H. ∎
Theorem 3 establishes that the adaptive learning rate of RMSProp, which scales updates based on past squared gradients, helps mitigate demographic parity gaps that can occur during training, whereas SGD does not offer such a mechanism. We relax the isotropic noise assumption in Theorem 3 and generalize the result to the anisotropic case in Appendix I.
@@ 4 Experiments
Setup. We use three public datasets that are widely used in this area: CelebA [31], FairFace [16], and MS-COCO [27]. CelebA and FairFace are datasets commonly used for group fairness analysis, while we also use a general-purpose dataset, MS-COCO, for further completeness. The details about these datasets are presented in Appendix J. The training protocol and backbones used in our experiments are described in Appendix K. We report the performance across different tasks using accuracy and F1. We evaluate group fairness using three widely recognized fairness criteria [38]: equalized odds, equal opportunity, and demographic parity. A predictor is fair under equalized odds definition if the true positive and false positive rates are equal across demographic subgroups. We measure the violation of this criterion as:
$F_{EOP}=\min_{i,j}(\frac{\sum_{c=1}^{C}\text{{TPR}}_{i,c}}{\sum_{c=1}^{C}\text{{TPR}}_{j,c}}),$ (27)
where $\text{{TPR}}_{i,c}$ denotes the true positive rate for the $i^{\text{th}}$ subgroup on class $c$. Finally, a predictor is said to be fair from a demographic parity point of view if the predicted label distribution is independent of sensitive attributes. The corresponding measure is defined as:
$F_{DPA}=\min_{i,j,q}\frac{p(\hat{y}=c_{i}|z_{j})}{p(\hat{y}=c_{i}|z_{q})}.$ (28)
For all the above fairness metrics, higher values indicate better (fairer) outcomes.
Implementation details. We use PyTorch [42] and up to four Nvidia A100 GPUs. The hyperparameter values are provided in Appendix L, which have been optimized using Weights & Biases [4].
@@
Regarding $F_{DPA}$, Adam consistently outperforms SGD, while RMSProp surpasses SGD in 4 out of 5 scenarios. Both optimizers achieve zero $F_{DPA}$ on MS-COCO with gender as the sensitive attribute. Interestingly, the largest discrepancy between RMSProp and SGD occurs on the FairFace dataset with age as the sensitive attribute. This aligns with our theoretical insights from Theorem 1. Specifically, FairFace exhibits extreme imbalance, with a minority-to-all ratio of approximately 0.9%, whereas CelebA with gender as the sensitive attribute has a substantially higher ratio of 42%. These findings highlight that as dataset imbalance intensifies, the gap between adaptive optimizers and SGD becomes more evident. We refer the reader to Appendix M for the results on fairness metrics for other backbones across CelebA, FairFace, and MS-COCO datasets with different sensitive attributes. The results are well-aligned with the results above. To further validate our findings across different variants of optimizers, we conduct additional experiments with SGD w/ momentum [30], AdamW [32], and AdaBound [33] optimizers. The results are provided in Appendix N.
Table 1: Fairness comparison across optimizers with tuning learning rate, decay rate, and batch size.
Gender Age
Adam RMSProp SGD Adam RMSProp SGD
$F_{\text{EOD}}$ 65.21 65.18 62.66 72.34 71.99 68.40
$F_{\text{EOP}}$ 99.90 99.91 96.60 99.10 99.10 97.77
$F_{\text{DPA}}$ 73.50 73.68 60.80 98.20 98.20 95.40
Beyond tuning the learning rate, we further investigate the influence of additional hyperparameters on fairness outcomes. Specifically, we extend the hyperparameter tuning using Bayesian Optimization over the learning rate, decay rate, and batch size. This comprehensive search is performed using the ViT backbone on the CelebA dataset, considering gender and age as sensitive attributes. The results are reported in Table 1. As shown, adaptive optimizers like Adam still achieve better fairness compared to SGD. This confirm that the fairness trends observed in our main study remain consistent even when multiple hyperparameters are jointly tuned.
@@
To demonstrate that our training is successful and the models are well-behaved, we evaluate the classification performance of ViT on the three datasets for facial expression recognition, multi-label classification, and gender classification, respectively. The results are presented in Table 2. The performance of the other models are presented in Appendix O. Here, we can confirm that in most cases, all three optimizers exhibit comparable accuracy and F1 scores across all tasks. For instance, on the CelebA dataset, SGD and RMSProp achieve an accuracy of 91%, while Adam attains 92%. This trend is consistent with previous findings in the prior work [34]. As evident, the improved fairness observed with adaptive gradient algorithms cannot be solely attributed to their performance. For instance, on the CelebA dataset, the F1 score for SGD is higher than that of RMSProp, whereas on FairFace, RMSProp achieves a higher F1 score than SGD. However, in both cases, RMSProp demonstrates better fairness performance, as shown in Fig. 2. This further supports our theoretical conclusion that RMSProp is fairer than SGD, regardless of overall performance.
@@
$F_{EOD}$ $\times 10^{-3} 1×1$$\times 10^{-3}$ $\times 10^{-3} 1×1$$\times 10^{-3}$
$F_{EOP}$ $\times 10^{-3} 1×1$$\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-3}$
@@
To rigorously assess the significance of fairness improvement of Adam and RMSProp over SGD, we conduct the following experiment. We train ViT on CelebA dataset 10 times using SGD, RMSProp, and Adam, while recording fairness metrics for both gender and age attributes in each run. To evaluate statistical significance, we perform a paired Wilcoxon signed-rank test [3] on the fairness metrics and report the corresponding p-values in Table 3. As shown in the table, most scenarios yield p-values in the order of 0.001, indicating a statistically significant improvement in fairness for adaptive gradient optimizers compared to SGD. The worst p-value in the table corresponds to SGD vs. RMSProp on CelebA with age as the sensitive attribute. However, even this value remains statistically significant under the conventional threshold of 0.05 [14, 40].
@@
Theorems 2 and 3 are general with respect to the loss function, which suggests their fairness benefits should be complementary to existing fairness-enhancing methods that add regularization terms to the primary loss. To validate this hypothesis, we conduct an experiment where we test the impact of optimizers in addition to the fairness-enhancing method [47] on two tabular datasets (ProPublica COMPAS [2] and AdultCensus [19]) with gender as the sensitive attribute. The results are presented in Table 4 for SGD, Adam, and RMSProp. Following [47], we report the average gap in equal opportunity, gap in equalized odds, and gap in demographic parity (lower is better) over 10 runs. Regarding the fairness enhancing method, across all metrics on both datasets, Adam and RMSProp achieve substantially lower (better) fairness gaps than SGD. This experiment demonstrates that the fairness benefits of adaptive optimizers are not limited to standard training but are complementary to and can amplify the effectiveness of existing fairness interventions. Additionally, Table 4 shows that in the absence of any fairness-enhancing method, adaptive optimizers yield better fairness outcomes than SGD. This finding further indicates that the fairness advantage of adaptive optimizers extends beyond computer vision applications.
Table 4: Comparison of fairness metrics across optimizers, with and without fairness-enhancing methods. Lower values indicate better fairness.
Gap in Equal Opportunity Gap in Equalized Odds Gap in Demographic Parity
Dataset Adam RMSProp SGD Adam RMSProp SGD Adam RMSProp SGD
With fairness-enhancing
ProPublica COMPAS 0.45 0.48 0.71 2.79 2.78 2.90 0.86 0.86 2.60
AdultCensus 3.15 3.16 3.38 2.39 2.41 4.28 7.00 6.80 11.79
Without fairness-enhancing
ProPublica COMPAS 13.99 13.90 15.19 13.99 13.95 14.98 11.49 11.45 11.80
AdultCensus 21.04 20.91 21.29 20.90 20.91 21.14 12.19 12.23 12.42
@@ 5 Conclusion
This paper provided both theoretical and empirical evidence that the choice of optimizer can significantly affect the group fairness of resulting deep learning models. Through establishing an analytically tractable setup and analyzing it via stochastic differential equations, we showed that SGD and RMSProp can converge to different minima with regards to fairness, with RMSProp more frequently reaching fairer optima. Next, we provided two theorems that showed 1) adaptive gradient methods (e.g., RMSProp) shrink subgroup-specific gradients more effectively, thereby reducing unfair updates compared to their SGD counterparts; 2) In a single optimization step, the worst case demographic disparity of RMSProp has the upper bound given by that of SGD. We validated these theoretical findings on three diverse datasets, namely CelebA, FairFace, and MS-COCO, and on multiple tasks. Across a range of group fairness criteria, i.e, equalized odds, equal opportunity, and demographic parity, adaptive optimizers consistently led to fairer solutions. We hope our work will encourage broader adoption of adaptive optimizers in fairness-critical domains, and we envision this research as a step toward developing more equitable deep learning models.
@@ Acknowledgment
This work was partially funded by the Natural Sciences and Engineering Research Council of Canada (NSERC), and was enabled in part by the support provided by the Digital Research Alliance of Canada.
@@ Appendix B Fokker-Planck equation
Also, suppose that $Q(x)$ is a function which is 0 outside of a bounded interval. For simplicity, we use $X$ instead of $X(t)$. By Itô’s lemma, we have:
@@
$\begin{split}\mathbb{E}[Q(X)-Q(0)]=\int_{0}^{t}\mathbb{E}[\left(M(\tau,X)Q^{\prime}(X)+0.5N^{2}(\tau,X)Q^{\prime\prime}(X)\right)]d\tau+\\ \int_{0}^{t}\mathbb{E}[N(\tau,X)Q^{\prime}(X)dB(\tau)].\end{split}$ (34)
@@
$\int_{-\infty}^{+\infty}\frac{\partial}{\partial t}p(t,x)Q(x)=\int_{-\infty}^{+\infty}p(t,x)M(t,x)Q^{\prime}(x)dx+\int_{-\infty}^{+\infty}p(t,x)0.5N^{2}(t,x)Q^{\prime\prime}(x)dx$ (36)
@@
$\int_{-\infty}^{+\infty}\left(\frac{\partial}{\partial t}p(t,x)+\frac{\partial}{\partial x}(p(t,x)M(t,x))-0.5\frac{\partial^{2}}{\partial x^{2}}(p(t,x)N^{2}(t,x))\right)Q(x)dx=0$ (37)
@@ Appendix C Well-behaved NGOS
Noisy Gradient Oracle with Scale Parameter (NGOS) models stochastic gradients as $g(w)=\nabla\mathcal{L}(w)+\Theta z$, where $\Theta$ is a noise scale parameter, and $z$ is a random variable that follows a distribution with covariance matrix $\Sigma(w)$ and a mean of zero. An NGOS is well-behaved if the conditions below are satisfied [37]:
@@
[…]
Gaussian noise is a well-behaved NGOS. However, the noise in NGOS is not restricted to Gaussian. Particularly, if the noise distribution is not heavy-tailed, then it satisfies requirements 4 and 5. Whether or not the noise distribution in real-world applications is heavy-tailed is still an open problem. Nevertheless, several experimental results denote that the conditions described above are not too strong [37].
@@ Appendix D Approximation Quality of SDE for SGD
The dynamics of SGD can be approximated by the following SDE [25]:
@@
where $W_{t}$ represents an approximation of $w_{k}$ at discrete time steps $k\,\eta$ and $\Sigma$ is the covariance matrix of the underlying NGOS. The noises that make the individual paths for the above SDE and the SGD are independent processes. Hence, SGD and its SDE approximation do not share the same sample paths; rather, they are close to each other in “distribution” or “weak sense” [25]. To formalize this mathematically, suppose $\Xi$ be the set of functions, $f(x)$, of polynomial growth, i.e. $\exists a,b\in\mathbb{R}^{+}$ such that $|f(x)|<a(1+|x|^{b})$. Then, SDE defined in Eq. 39 is the order $r$ weak approximation of SGD if for each function $f(x)$ in $\Xi$, there exists a constant $C$ such that for all $k=1\dots N$, $|\mathbb{E}[f(W_{k\eta})]-\mathbb{E}[f(w_{k})]|<C\eta^{r}$. This is a well-known result in analyzing optimization algorithms using SDEs, first time proposed in [25]. Similar theorems for RMSProp and Adam are available in [37].